# Adaptive Parametric Activation

### v2

@@ 5 Results
Generalisation to Balanced tasks. In Table 6-a, we show that APA*+AGLU increases the performance of SE-MaskRCNN-R50 by $0.7$pp on $AP^{b}$ and $AP^{m}$ on COCO. In Table 6-b, we show that APA*+AGLU increase the performance of SE-FasterRCNN-R50 by $2.9$pp and the performance of SE-Cascade-RCNN-R50 by $2.1$pp on the challenging V3Det, which are significant increases considering that this dataset has 13K classes. In Table 6-c, we show that our methods increase the performance for R50, SE-R50 and CBAM-R50, by $0.6$pp, $1.2$pp and $0.6$pp respectively. Finally in Table  6-d, we show that our modules increase the performance by $1.2$pp on SE-R50, $0.9$pp on SE-R101 and $0.5$pp on SE-R101, showing that they can generalise in balanced training.
@@
Table 6: Resuls on balanced tasks. In (a) and (b), we combine APA*+AGLU on COCO dataset and V3Det dataset respectively. In (c) we combine AGLU and APA* to ResNet50, SE-ResNet50 and CBAM-ResNet50 in ImageNet1K. In (d) we show that APA*+AGLU generalises to deeper backbones in ImageNet1K.
@@
FasterRCNN-R50 25.4
@@
APA*+AGLU-CRCNN 35.4
Method top-1
ResNet50 [37] 76.9
w/ AGLU 77.5
SE-ResNet50 [37] 77.5
w/ APA*+AGLU 78.7
CBAM-ResNet50 [37] 78.3
w/APA*+AGLU 78.9
Method top-1
SE-R50 77.5
+APA* 77.9
+APA* + AGLU 78.7
SE-R101 79.4
+APA* 79.2
+APA* + AGLU 80.3
SE-R152 80.3
+APA* 80.5
+APA* + AGLU 80.8
@@
(b)V3Det
(c)ImageNet1k
(d) ImageNet1k
6 Related Work
Long-tailed image recognition. Long-tailed image recognition can be grouped according to representation learning and classifier learning techniques. Representation learning techniques improve the feature quality through rare class features generation [85, 87, 104], contrastive objectives [91, 15, 117, 74, 53, 43], ensemble or fusion models [95, 51, 14, 113, 50, 13, 7], knowledge distillation [51, 41, 29, 52], knowledge transfer [58, 67, 118] and data augmentation [111, 68, 101]. Classifier learning techniques enhance rare class classification through decoupled training [44, 108, 92, 47, 36], margin adjustment [63, 71, 34, 8, 39, 65, 89, 93, 110, 2], cost-sensitive learning [16, 46, 96, 38], resampling [9, 61, 76, 120, 22, 68, 33, 104], dynamic loss adaptation based on batch statistics [35, 80, 94], gradient statistics [79, 49], weight norms [93] and classification scores [20, 30]. These techniques are efficient, however they are applied either during the input phase or during the loss disregarding the intermediate channel activations.
Attention Networks. Attention networks, such as spatial attention [42], channel attention [37] and spatial-channel attention [98], are widely used in image recognition. Self-attention is the core mechanism of the Visual Transformer (ViT) architecture [19], and recently many improvements have been made to the Transformer architecture [81, 83, 56] and its training procedure [78, 6, 82].
In our work, we primarily used APA with channel attention models, but it also works with plain ResNets, Spatial and Spatial-Channel attention models. In the Appendix, we show results with self-attention models.
Activation functions The RELU function [21] dominates the landscape of deep image classification, and it is especially used inside the convolutional networks, while the GELU and SiLU [31] are more commonly used inside the transformer network [19] or the ConNext models [57, 97]. The PRELU function [27] is a generalisation of the RELU, because it linearly activates both positive and negative inputs and the ELU [11] is a follow-up work that enforces saturation on the negative inputs after some threshold. Our AGLU activation, generalises the RELU, the GELU and SiLU and it can also be seen as a smoother version of PRELU.
7 Conclusion
@@
Our work highlights the impact of the activation function inside the model’s activations for balanced and imbalanced data distributions. We have empirically showed that the degree of data imbalance affects the logit distributions and the intermediate signals and we have showed that the commonly used Sigmoid activation function is unable to model the intermediate features. To this end, we have proposed an novel adaptive parametric activation that unifies most common activation functions under the same formula and we have tested it in several long-tail and balanced classification and detection benchmarks showing great generalisation.
Acknowledgements. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) project “ViTac: Visual-Tactile Synergy for Handling Flexible Materials” (EP/T033517/2).
Appendix 0.A Representations’ Quality
@@
We evaluate the quality of the representations learned by the SE and APA* models using the recently proposed Neural Collapse framework [66].
@@
$\begin{split}\Sigma_{W}:=&\frac{1}{n}\sum_{k=1}^{K}\sum_{j=1}^{n_{k}}(f_{k,j}-\bar{f_{k}})(f_{k,j}-\bar{f_{k}})^{\top}\\ \Sigma_{b}:=&\frac{1}{K}\sum_{k=1}^{K}(\bar{f_{k}}-f_{G})(\bar{f_{k}}-f_{G})^{\top}\\ \end{split}$ (11)
The $\Sigma_{W}$ matrix shows how distant are individual features $f_{k,j}$ from their class prototype $\bar{f_{k}}$ and it is an indicator of feature compactness. The $\Sigma_{b}$ matrix shows how distant are the class prototypes from the global feature, indicating the class separability. Using these matrices we measure the Neural collapse Variability $NC1$ according to [119] as follows:
@@
where the $\dagger$ symbol denotes the pseudo inverse of $\Sigma_{b}$. $NC1$ measures the magnitude of the within-class covariance $\Sigma_{W}$ compared to the magnitude of the between-class covariance $\Sigma_{b}$ as explained in [119].
In practise, a low $NC1$ measure shows that the model has more compact features since $\Sigma_{W}\downarrow$ decreases and more separable class prototypes because the $\Sigma_{b}\uparrow$ increases. Having more compact features and more separable class prototypes make the representations better and enhance the classification results as shown empirically in previous works [112, 102, 103, 99].
Using Equation 12, we measure the $NC1$ of the deep features of the penultimate layer of SE and APA*, in Table 7 using ImageNet-LT test-set. As the results suggest, our APA* has lower $NC1$ measure for all backbones, showing that APA* produces superior representations that are more compact and seperable than the baseline. This provides another qualitative explanation why our APA* has better performance than SE.
Table 7: Neural Collapse NC1 measure, on ImageNet-LT test set. APA* has lower $NC1$ measure than the baseline, which indicates that it has learned superior representations.
@@
Appendix 0.B Implementation Details
The implementation details of APA* and AGLU are shown in Table 8. For balanced ImageNet-1K, the $\lambda$ parameters are initialised as random variables drawn from a Uniform distribution (U), with low parameter 00, and high parameter $1.0$. The APA $\kappa$ parameters are initialised with $U(0,1)$ and the AGLU $\kappa$ parameters are with initialised with $U(0.8,1.2)$. For all other downstream tasks, that use a pretrained model, such as COCO, LVIS, Places-LT and V3Det, we don’t re-initialise the $\kappa$ and $\lambda$ parameters and we simply load them from the pretained ImageNet1K model.
0.B.1 Stable APA implementation
@@
CutMix $\alpha$ 1.0 1.0 - -
Label smoothing $\epsilon$ 0.1 0.1 - -
@@
AGLU $\kappa$ Init U(0.8,1.2) U(0.8,1.2) N/A U(0.8,1.2) N/A
@@
Table 8: Implementation details for Long-tailed Datasets, across various architectures.
0.B.2 AGLU derivatives
@@
$=x^{2}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{-1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
@@
$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial x}$ $=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$ (16)
$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\cdot\partial\frac{(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\kappa\lambda\exp(-\kappa x))$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
$=\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}$
Appendix 0.C Results
0.C.1 Experiments with AGLU and plain ResNets
In Table 9, we show the result of AGLU when it applied inside plain ResNet50 models, (i.e. without channel attention). As the Table suggests, by simply replancing the RELU with AGLU, our method consistently increases the performance of plain ResNet models.
Table 9: Top-1 accuracy on long-tailed classification datasets using ResNets.
Dataset CIFAR100-LT ImageNet-LT iNaturalist
Imbalance factor 10 100 256 500
Model ResNet-32 ResNet50 ResNeXt50 ResNet50
RELU 65.7 51.8 55.0 57.0 69.9
AGLU (ours) 66.8 52.0 56.0 57.6 72.4
0.C.2 Experiments with AGLU and Vision Transformers on ImageNet1K
We perform a preliminary experiment with Vision Transformer models such as ViT [19] and Swin [56] using ImageNet1K. We replace the GELU activation with AGLU in every feedforward layer and we keep all other settings the same. As shown in Table 10, AGLU performs comparably to GELU. We believe this is because the Self-Attention function makes the features smooth, by removing their harmonising components and it makes them more Gaussian-like [90, 114, 18, 5]. Consequently, the Gaussial linear error unit, GELU, might be a good choice for the ViT network and our AGLU method has comparable performance.
Table 10: Results of AGLU using ViT models on ImageNet1K.
Model Activation epochs top-1
ViT-B [19] GELU 200 78.3
ViT-B [19] AGLU 200 78.5
Swin-T [56] GELU 100 78.7
Swin-T [56] AGLU 100 78.9
0.C.3 Impact of Initialisation
In all of our experiments, we have initialised $\lambda$ using the Uniform distribution with low parameter 0 and high parameter 1 as a default. Regarding the $\kappa$ parameter inside AGLU, we initialise it to be close to $1.0$, as this works best, as shown in Table 11. Regarding the $\kappa$ parameter inside the attention layer, we found that initialising it with $Uniform(-1,0)$ is slightly better than $Uniform(0,1.0)$ as shown in Table 12.
Table 11: AGLU-$\kappa$ parameter initialisation, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.

### v3

@@ VI-A Image Classification
In Table IX-a, we show that APA*+AGLU increases the COCO performance of SE-MaskRCNN-R50 by 0.70.7pp on $AP^{b}$ and $AP^{m}$. Also, it increases the performance of SE-MaskRCNN-R101 by 0.90.9pp on $AP^{b}$ and 0.50.5pp on $AP^{m}$. As Table IX-b shows, APA*+AGLU improves $AP^{b}$ by 1.11.1pp and 1.61.6pp on the vast Objects365v2 dataset using the SE-FasterRCNN-R50 and R101 models respectively. In Table IX-c, we show that APA*+AGLU increase the performance of SE-FasterRCNN-R50 by 2.92.9pp and the performance of SE-Cascade-RCNN-R50 by 2.12.1pp on the challenging V3Det, which are significant increases considering that this dataset has 13K classes.
@@
TABLE IX: Resuls on balanced tasks. In (a) and (b) and (c), we combine APA*+AGLU on $CO$, Objects365v2 and V3Det dataset respectively. In (d) we combine AGLU and APA* to ResNet, SE-ResNet and CBAM-ResNet in ImageNet1K. The shows that the baseline is taken from mmdetection [mmdetection].
@@
FasterRCNN-ResNet50 19.8
SE-FasterRCNN-ResNet50 26.4
w/ APA*+AGLU 27.5
SE-FasterRCNN-ResNet101 29.3
w/ APA*+AGLU 30.9
Method $AP^{b}$
FasterRCNN-Resnet50 25.4
@@
w/ APA*+AGLU 35.4
@@
(b) Objects365v2
(c) V3Det
VI-C Experiments with Vision Language Model.
We further test the performance of our proposed activation function in the visual instruction finetuning task using vision-language models (VLMs), following the LLAVA series paradigm [li2024llava, liu2023llava, liu2024llavanext, li2024llavanext-ablations, li2024llavanext-strong]. In more detail, we use an MLP with AGLU activation as a connector module and pretrain two VLMs, including the Qwen2 [yang2024qwen2technicalreport] and Deepstack [meng2024deepstack] with the SigLIP-so400M [zhai2023sigmoid] vision encoder, using the LLaVA-1.5-PT dataset that has 558k image-text pairs [liu2023improvedllava], keeping all parameters frozen except for the connector MLP. Next, we perform instruction tuning and we fully finetune the models using the OpenLLaVA-Next dataset [chen2024open]. To evaluate the performance, we use the LMM-evaluation kit [zhang2024lmmsevalrealitycheckevaluation] that includes many vision-language benchmarks such as MME [fu2023mme], AI2D [kembhavi2016diagram], DocVQA [mathew2021docvqa], GQA [hudson2019gqa], MMB [liu2024mmbench], OCRB [Liu_2024_ocrbench], MMMU [yue2024mmmu], OKVQA [marino2019ok], POPE [li2023evaluating], SQA [lu2022learn], TextVQA [singh2019towards] and VisWiz [gurari2018vizwiz]. As shown in Table X, our AGLU connector increases the performance, in both Qwen2-1.5B, Deepstack-7B and Qwen2-7B models by 2.42.4pp, 0.60.6pp and 0.90.9pp, respectively, on the macro average metric and achieves superior performance compared to open source models like LLaVA-Next-7B [liu2024llavanext] and proprietary models like DeepSeek-Janus-Pro 7B. This indicates that AGLU is particularly useful in vision language training because its adaptive behaviour can bridge the domain gap between language and image distributions.
TABLE X: Comparative results on the visual instruction following task, using a Qwen2-SigLIP model.
[…]
DeepSeek-Janus-Pro-7B [chen2025janus] 1567 - - 59.1 - - 79.2 - 41.0 - - 87.4 - - - -
LLaVA-1.5-7B [liu2023improvedllava] 1511 - - 62.0 - 58.3 64.3 - - - - 86.0 71.6 61.3 50.0 -
LLaVA-Next-7B [liu2024llavanext] 1519 66.6 74.4 64.2 - 60.6 67.4 - 35.8 - - 86.5 70.1 64.9 57.6 -
VILA-7B [lin2024vila] 1533 - - 62.3 - 61.7 70.3 - - - - 85.5 68.3 64.4 57.8 -
InternVL-Chat-7B [chen2024internvl] 1525 - - 62.9 - - - - - - - 87.1 - 57.0 52.5 -
Qwen2-1.5B-SigLIP 1393 59.3 46.6 60.3 33.7 57.3 57.6 33.8 27.9 42.5 51.4 86.5 70.3 53.4 37.5 52.5
w/ AGLU 1390 60.3 47.7 60.9 37.7 60.1 64.3 47.6 26 41.3 52.3 87.5 69.8 54.9 43.4 54.9
Deepstack-7B [meng2024deepstack] 1592 74.6 73.2 64.4 46.7 72.8 75.3 51.0 47.1 53.7 56.2 88.0 71.4 66.1 61.7 65.5
w/ AGLU 1573 75.3 73.6 64.3 47.3 73.6 76.3 51.3 49.0 53.0 58.0 87.9 76.1 66.4 61.0 66.1
Qwen2-7B-SigLIP 1596 73.9 74.8 63.9 42.9 72.3 74.2 56.3 45.8 55.9 61.2 87.7 70.6 68.0 53.8 65.4
w/ AGLU 1634 74.1 75.6 64.3 46.9 71.8 75.1 55.1 45.7 58.3 61.9 88.1 76.5 68.0 51.0 66.3
VI-D Experiments in Image Generation.
We try our AGLU activation in the task of image generation. Specifically, we select LightningDIT-XL [yao2025reconstruction], which is a strong transformer-based, latent diffusion model as our baseline, and we replace all its activation functions with AGLU, keeping any other training details and hyperparameters fixed. We train the AGLU-LightiningDIT-XL model with patch size one, on the Imagenet1k-256 generation benchmark for 64 epochs and evaluate the generation quality using the ADM evaluation suite [dhariwal2021diffusion]. As Table XI shows our activation function achieves good performance, decreasing the FID score by $0.1$ and increasing the Recall by 0.80.8pp compared to LightningDiT baseline.
TABLE XI: AGLU achieves good performance in image generation.
Model FID sFID Inception Score Precision Recall
LightiningDIT-XL[yao2025reconstruction] 2.2 4.3 251.9 81.5 58.3
LightiningDIT-XL w/ AGLU 2.1 4.2 251.1 81.0 59.1
VI-E Experiments with Language Model.
We perform a language next-token prediction experiment using GPT2 [radford2019language] and the FineWeb-Edu [lozhkov2024fineweb-edu] subset that contains 10 billion GPT2 tokens. The model implementation follows nanogpt by [Karpathy2024]. We train four GPT2 models, GPT2-small that has 124 million parameters, GPT2-medium that has 350 million, GPT-large that has 774 million parameters and GPT-xl that has 1.5 billion parameters for 10 billion tokens, with a batch size of 0.5 million tokens, learning rate 6e46e-4 and Adam optimizer with momentum [kingma2014adam]. We test the model on the HellaSwag benchmark [zellers2019hellaswag] using zero-shot evaluation. To apply AGLU with GPT2, we simply replace all GELU activations with AGLU inside all MLP layers of the transformer. As the results suggest in Table XII, AGLU increases the performance of GPT2 by 0.2pp, 0.4pp, 0.3pp and 0.8pp on the HellaSwag benchmark using GPT2-small, medium, large and xl respectively, and reduces the validation loss on the FineWeb-Edu by 0.01 and 0.02 on GPT2-medium, and GPT2-large and xl respectively showing that AGLU is a good alternative for the next token prediction task, especially for deeper models.
TABLE XII: Comparative results using GPT2 model on the HellaSwag benchmark and Fineweb-Edu validation set.
Method HellaSwag FineWeb Val-Loss
GPT2-small 30.8 3.08
GPT2-small with AGLU 31.0 3.08
GPT2-medium 35.2 2.86
GPT2-medium with AGLU 35.6 2.85
GPT2-large 38.8 2.72
GPT2-large with AGLU 39.1 2.70
GPT2-xl 42.7 2.62
GPT2-xl with AGLU 43.5 2.60
VII Analysis
VII-A Computation cost.
AGLU costs more computing resources compared to non-parametric activation functions because it requires training. The additional computation depends on the architecture, the optimizer and the size of the layer where AGLU is applied. In Table XIII, we compare AGLU against the vanilla architecture for different models, in terms of the number of parameters, peak memory measured in MB, throughput measured in samples per second and FLOPS using ImageNet1K. AGLU requires $17\%$ more memory, it has $3\%$ smaller throughput and it has the same number of parameters and FLOPS in the SE-ResNet50 architecture. In the ViT-B architecture, it lowers the throughput by $3\%$ but keeps all other measures the same. The extra computing occurs mostly during training and it depends on the network’s architecture, i.e., if AGLU is applied after a small linear layer, then the training cost is low, however, if it is applied after a large layer, i.e. a large convolution layer, then the training cost is larger.
VII-B Learned activations.
We visualise the converged AGLU activation functions, i.e. the output $AGLU(z,\lambda,\kappa)=y$ for the ImageNet21K-ViT-B and GPT models. To better understand their behaviour, we measure AGLU’s expected output $E[y]$, for random inputs $z$ drawn from the Normal distribution. It is striking to notice that both the ViT and GPT2, have the same activation patterns, i.e. the activation of the early layers is harder, i.e., more non-linear and it allows mostly positive inputs to pass whereas in the deeper layers it becomes smoother, i.e. more linear, allowing both positive and negative outputs to propagate. We hypothesize that, given the models’ hierarchical feature extraction ability, the network distinguishes the early features easily and draw harder decision boundaries with the activation function, but as the layers grow deeper, the features become more abstract and the network distinguishes the features less easily and uses smoother decision boundaries.
Figure 7: Comparison of the expected AGLU-outputs $E[y]$ for ViT and GPT models, given input drawn from the Normal distribution, across layers. The expected value is larger at the first layers, indicating harder activations, and it becomes smaller at the last layers, indicating smoother activations, for both GPT and ViT.
TABLE XIII: Computation cost of AGLU. FLOPS are measured on a V100 GPU using batch-size of 1, $224^{2}$ image resolution and mixed precision. Peak memory (MB) and throughput ($\frac{\text{img}}{\text{sec}}$) are measured on the ImageNet1K finetuning.
Model Parameters MB $\frac{\text{img}}{\text{sec}}$ GFLOPS
SE-R50 28M 6020 393.6 8.2
w/ APA*+AGLU 28M 7064 382.7 8.2
ViT-B 86M 5145 263.7 33.5
w/ AGLU 86M 5145 256.9 33.5
VIII Discussion
We have shown that the adaptive activation function is a solid alternative to standard activation functions such as RELU or GELU, for many vision tasks, especially for long-tailed tasks. Some of the biggest benefits are that, it is a generalisation of many common activation functions, it can be computed analytically and it is a easy to use. We believe that AGLU, can be particularly effective when applied to specialised modules such as in the VLM connector, or CNN detectors and is a good alternative for large scale pretrained Transformer models, showing 0.30.3pp increase in ViT and GPT models for both classification and generation tasks. This is lower than the performance increase observed with CNNs, potentially because the transformer is a highly non-linear structure and already very expressive due to its attention mechanism. Regarding, the computational resources, AGLU is overall more expensive than non-adaptive functions but the computing cost is diminishing when AGLU is applied to small linear layers.
IX Conclusion
@@
Our work highlights the impact of the activation function inside the model’s activations for balanced and imbalanced data distributions. We have empirically shown that the degree of data imbalance affects the logit distributions and the intermediate signals and we have shown that the commonly used Sigmoid activation function is unable to model the intermediate features. To this end, we have proposed an novel adaptive parametric activation that unifies most common activation functions under the same formula and we have tested it in several long-tail and balanced benchmarks showing great generalisation.
Appendix A Representations’ Quality
@@
We evaluate the quality of the representations learned by the SE and APA* models using the recently proposed Neural Collapse framework [papyan2020prevalence].
@@
[…]
The $\Sigma_{W}$ matrix shows how distant are individual features $f_{k,j}$ from their class prototype $\bar{f_{k}}$ and it is an indicator of feature compactness. The $\Sigma_{b}$ matrix shows how distant are the class prototypes from the global feature, indicating the class separability. Using these matrices we measure the Neural collapse Variability $NC1$ according to [zhu2021geometric] as follows:
@@
where the $\dagger$ symbol denotes the pseudo inverse of $\Sigma_{b}$. $NC1$ measures the magnitude of the within-class covariance $\Sigma_{W}$ compared to the magnitude of the between-class covariance $\Sigma_{b}$ as explained in [zhu2021geometric].
In practise, a low $NC1$ measure shows that the model has more compact features since $\Sigma_{W}\downarrow$ decreases and more separable class prototypes because the $\Sigma_{b}\uparrow$ increases. Having more compact features and more separable class prototypes make the representations better and enhance the classification results as shown empirically in previous works [Zhong_2023_CVPR, yang2022inducing, yang2023neural, xie2023neural].
Using Equation 12, we measure the $NC1$ of the deep features of the penultimate layer of SE and APA*, in Table XIV using ImageNet-LT test-set. As the results suggest, our APA* has lower $NC1$ measure for all backbones, showing that APA* produces superior representations that are more compact and seperable than the baseline. This provides another qualitative explanation why our APA* has better performance than SE.
TABLE XIV: Neural Collapse NC1 measure, on ImageNet-LT test set. APA* has lower $NC1$ measure than the baseline, which indicates that it has learned superior representations.
@@
Appendix B Implementation Details
The implementation details of APA* and AGLU are shown in Table XV. For balanced ImageNet-1K, the $\lambda$ parameters are initialised as random variables drawn from a Uniform distribution (U), with low parameter 0, and high parameter $1.0$. The APA $\kappa$ parameters are initialised with $U(0,1)$ and the AGLU $\kappa$ parameters are with initialised with $U(1,1.3)$. For all other downstream tasks, that use a pretrained model, such as COCO, LVIS, Places-LT and V3Det, we don’t re-initialise the $\kappa$ and $\lambda$ parameters and we simply load them from the pretained ImageNet1K model. We do not use weight decay for the $\kappa$ and $\lambda$ parameters as in [he2015delving].
B-A Stable APA implementation
@@
TABLE XV: Implementation details for Long-tailed Datasets, across various architectures.
@@
CutMix $\alpha$ - 1.0 1.0 - -
Label smoothing $\epsilon$ - 0.1 0.1 - -
@@
AGLU $\kappa$ Init U(1,1.3) U(1,1.3) N/A U(1,1.3) N/A
@@
B-B AGLU derivatives
@@
$=x^{2}\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda})}}{\lambda\exp(-\kappa x)+1}$
@@
$\frac{\partial AGLU(x,\kappa,\lambda)}{\partial x}=\partial\frac{x\cdot(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+x\cdot\partial\frac{(\lambda\exp(-\kappa x)+1)^{\frac{1}{-\lambda}}}{\partial x}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+x\frac{(\lambda\exp(-\kappa x)+1)^{(\frac{-1}{\lambda}-1)}}{-\lambda}\cdot(-\kappa\lambda\exp(-\kappa x))\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\exp(-\kappa x)\frac{(\lambda\exp(-\kappa x)+1)^{(-\frac{1}{\lambda})}}{\lambda\exp(-\kappa x)+1}\\ =\eta_{\text{ad}}(x,\lambda,\kappa)+\kappa x\frac{\eta_{\text{ad}}(x,\lambda,\kappa)}{\lambda+\exp(\kappa x)}\\$ (16)
Appendix C Results
C-A Impact of Initialisation
In all long-tail experiments, we have initialised $\lambda$ using the Uniform distribution with low parameter 0 and high parameter 1 as a default. Regarding the $\kappa$ parameter inside AGLU, we initialise it to be close to $1.0$, as this works best, as shown in Table XVI. Regarding the $\kappa$ parameter inside the attention layer, we found that initialising it with Uniform(1,0)Uniform(-1,0) is slightly better than $Uniform(0,1.0)$ as shown in Table XVII. In all other experiments including image generation, ImageNet21K, ImageNet1K training, VLM training and GPT training we used $\kappa$ and $\lambda$ initialization from the $Uniform(0.8,1.2)$ distribution, to simplify the codebase. The latter initialization has good performance across many balanced tasks. We note that it might be possible to achieve better performance by tuning the initialisation for every task, however we chose not to do it due to computing constrains.
TABLE XVI: AGLU-$\kappa$ parameter initialisation, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.